# Permutation twisted cohomology, remixed

### v2

@@ 6. Injectivity of the comparison map
We informally say some words about noetherianity to close, to hopefully conjure hope within the reader that for any finite $p$-group $G$, $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is noetherian. First, one can show $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ is noetherian when $G$ satisfies the following property: the indecomposable representative of each irreducible endotrivial $C$ in $\mathcal{B}(G)$ has at most one indecomposable module in every homological degree. This holds for instance (and possibly only when) $G$ is abelian, where the endotrivials correspond to subgroups $N\leq G$ such that $G/N$ is cyclic (see [Mil24, Section 6]). With this property, the proof follows essentially identically as [BG25, Lemma 12.12], with that the maps $\iota^{1}_{C}$ and $\iota^{G}_{C}$ replacing the maps $b_{N}$ and $a_{N}$ respectively for each irreducible endotrivial $C\in\mathcal{B}(G)$.
One crux of the matter is that if an irreducible endotrivial has a homological degree with two or more indecomposables, the exact sequence [BG25, 12.14] sends a term $f\in\operatorname{Hom}_{\mathscr{K}(G)}(k,v\otimes u^{\otimes q})$ to a direct sum of morphisms $(f_{1},\dots,f_{n})$ with $f_{i}\in\operatorname{Hom}_{\mathscr{K}(H_{i})}(k,\operatorname{Res}^{G}_{H_{i}}(v)[q])$, where the subgroups $H_{i}\leq G$ are the corresponding stabilizers of the indecomposable permutation modules. It is unclear if one can inductively lift this sum of morphisms to a morphism in $\operatorname{Hom}_{\mathscr{K}(G)}(1,v\otimes u^{\otimes q})$, polynomial in morphisms $k\to C[s]$ with $C\in\mathcal{B}(G)$, such that its image in every $\operatorname{Hom}_{\mathscr{K}(H_{i})}(k,\operatorname{Res}^{G}_{H_{i}}(v)[q])$ corresponds to the image of $f$. Additionally, the lengths of the irreducible endotrivials $C$ can be arbitrarily long, but an inductive argument may be used to resolve this issue.
A potential roadmap to proving noetherianity modulo nilpotents is to verify that the image of the map $\prod\hat{\Psi}:\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\prod_{H\in s_{p}(G)/G}\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H})$ is finite, as the kernel of $\prod\hat{\Psi}$ is nilpotent (Proposition 5.6). One can further reduce this to the elementary abelian case by postcomposing by restriction to all elementary abelian subquotients of $G$, as the restriction $\operatorname{H}^{\scriptscriptstyle\bullet}(G)\to\prod_{E\in\operatorname{elemab}}(G)\operatorname{H}^{\scriptscriptstyle\bullet}(E)$ detects nilpotents as well (see [Ben98b, Proposition 5.2.2]). It is not too unreasonable to expect this to hold - for instance [Eve61, Theorem 7.1] states that for any subgroup $H\leq G$, $\operatorname{H}^{\scriptscriptstyle\bullet}(H)$ is finite in $\operatorname{H}^{\scriptscriptstyle\bullet}(G)$ via restriction. Hence $\operatorname{H}^{\scriptscriptstyle\bullet}(H)^{G}$ (i.e. the image of restriction, see [Ben98a, Proposition 3.8.2]), is finite in $\operatorname{H}^{\scriptscriptstyle\bullet}(G)$. For this reason, it is quite pertinent and interesting to determine the image of $\hat{\Psi}^{H}:\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)\to\operatorname{H}^{\scriptscriptstyle\bullet}({G}/\!\!/{H})$ for any subgroup $H\leq G$. We note for instance that one can show the image of $\prod\hat{\Psi}$ is finite when $p=2$ and $G=Q_{8}$ exploiting the fact that every subgroup of $Q_{8}$ is normal. However, it is not clear if an analogue of [Bal10a, Theorem 7.3] under the hypothesis that the cohomology ring is noetherian modulo nilpotents can be utilized.

### v3

@@ 7.1. Noetherianity for groups of order $p^{3}$
The identification $\ker(\operatorname{Res}^{G}_{H})=\operatorname{im}(t^{G}_{H})$ does some heavy lifting in the two proofs: it is a seemingly necessary condition to perform the ‘copying’ argument. This does not hold in full generality for subgroups of a $p$-group for morphisms of chain complexes of index at least $p^{2}$; simple counterexamples are easy to construct. However, if one considers the irreducible endotrivials $C\in\mathcal{B}(D_{2^{n}})$ with $n\geq 3$ (in particular, the three-term complexes), surprisingly, the same equality can be shown to hold for all subgroups, not just those of index 2. This seems to be a nice quirk of working up to homotopy. It is not clear to us if such an equality holds for all effective endotrivials for arbitrary groups, or if such an equality is necessary for Noetherianity of $\operatorname{H}^{{\scriptscriptstyle\bullet}{\scriptscriptstyle\bullet}}(G)$ to hold.
More generally, the proposition we’d like to show is as follows. Let $C$ be an effective endotrivial with $n+1$ terms, let $d_{i}$ be a transition map for $0<i\leq n$, let $\hat{C}$ be formed from $C$ by truncation with $\hat{C}_{j}=0$ for $j>i$, and let $\phi$ denote the obvious morphism $\hat{C}\to C_{i}[i]$. We write $C_{i}=\bigoplus_{l=1}^{m}k[G/H_{l}]$. Then we expect the induced homomorphism
$\operatorname{Hom}_{\mathscr{K}(G)}(k,D\otimes\hat{C})\xrightarrow{\phi^{*}}\operatorname{Hom}_{\mathscr{K}(G)}(k,D\otimes C_{i}[i])\cong\bigoplus_{l=1}^{m}\operatorname{Hom}_{\mathscr{K}(H_{l})}(k,\operatorname{Res}^{G}_{H_{l}}D[i])$
has image consisting of $G$-stable elements. Such a result need not occur for all complexes arising from the orbit category; this appears to be a special property of effective endotrivials.
If indeed this holds, a combination of the recursive argument in Theorem 7.4 on the length of the endotrivial $C$, the ‘copying’ argument in Theorem 7.8 relying on $G$-stability, and Hilbert’s basis theorem to construct a finite $G$-stable generating set, should suffice to prove Noetherianity. We leave the full argument to the reader. Unfortunately, despite the progress we have made elucidating the structure of effective endotrivials, such a result at present still appears far out of reach.